# Spin excitations of the Shastry-Sutherland model – altermagnetism and proximate deconfined quantum criticality

### v1

@@ Abstract
Acknowledgments.— We thank Runze Chi, Wenan Guo, Bruce Normand, Wei Li, Zhengxin Liu, Anders W. Sandvik, Yiming Wang, Jize Zhao for helpful discussions. This work is supported by the National Key R&D Program of China (Grant No. 2023YFA1406500), the National Natural Science Foundation of China (Grant Nos. 12334008, 12274458, 12174441).
SUPPLEMENTAL MATERIAL – Spin excitations of the Shastry-Sutherland model – altermagnetism and proximate deconfined quantum criticality
Hongyu Chen1, Guijing Duan1, Changle Liu2, Yi Cui1,3, Weiqiang Yu1,3, Z.Y.Xie1,3, Rong Yu1,3
1School of Physics and Beijing Key Laboratory of Opto-electronic Functional Materials and Micro-nano Devices, Renmin University of China, Beijing 100872, China
2School of Engineering, Dali University, Dali, Yunnan 671003, China
3Key Laboratory of Quantum State Construction and Manipulation (Ministry of Education), Renmin University of China, Beijing, 100872, China
@@
I iPEPS method for calculating the ground state and excited states
@@
Figure S-1: (a): The iPEPS setup in the calculation of the ground state of the SS model. The tensors are arranged at the sites of the SS lattice. The bright-colored sites indicate a $2\times 2$ unit cell used in the calculation. (b): Illustration of the single-mode iPEPS ansatz for the calculation of excited states. In each term of the summation, a single local tensor $A_{\alpha}$ located at site $r$ belonging to a specified sublattice is modified to a different tensor $B_{\alpha}$. Here $\alpha=1,2,3,4$, corresponding to a $2\times 2$ unit cell.
@@
In this work, we apply the infinite projected entangled pair state (iPEPS) method [56, 57] to calculate the ground state of the Shastry-Sutherland (SS) model. A sketched iPEPS setup with a $2\times 2$ unit cell is shown in Fig. S-1(a). In the iPEPS setup, a local tensor is defined on each site of the SS lattice. As shown in Fig.  S-1, each local tensor has a physical index corresponding to the spin degrees of freedom, with dimension $2$ and denoted as curve, and four virtual indices corresponding to the four links of the site, with dimension $D$ referred to as the iPEPS bond dimension and denoted as straight. In this work, as emphasized as bright color in Fig. S-1, the unit cell is chosen to be $2\times 2$, and this is the minimal size compatible with the ground state of the SS model.
@@
$E=\frac{\sum_{\langle ij\rangle}\langle\Psi|\hat{H}_{ij}(J)|\Psi\rangle}{\langle\Psi|\Psi\rangle}+\frac{\sum_{\langle\langle ij\rangle\rangle}\langle\Psi|\hat{H}_{ij}(J^{\prime})|\Psi\rangle}{\langle\Psi|\Psi\rangle}$ (S-1)
where $\langle ij\rangle$ and $\langle\langle ij\rangle\rangle$ denote nearest neighbors and second nearest neighbors, respectively. In this sense, the energy can be regarded as a complicated function of the local tensors, the variational parameters in the iPEPS ansatz. The denominators and numerators in Eq. (S-1) can be effectively calculated by using the corner transfer matrix renormalization group method [58, 59, 60, 75]. Then, starting from an arbitrary state, the ground state is obtained by variationally minimizing the energy through gradient-based optimization, e.g., the L-BFGS optimization strategy [76]. In this work, the energy gradients concerning the local tensors are obtained effectively by the automatic differentiation techniques [61, 62].
Once the ground state $|\Psi_{0}\rangle$ is obtained, we can construct the iPEPS ansatz for the excited state through the single-mode approximation [63, 64, 65]. More specifically, the excited ansatz with lattice momentum $k$ reads as
$|\Psi_{k}(B_{\alpha})\rangle=\sum_{r\in\alpha}e^{ikr}|\Psi_{r}(B_{\alpha})\rangle$ (S-2)
@@
$\langle\Psi_{r}(B_{\alpha})|\Psi_{0}\rangle=0.$ (S-3)
In practice, $B_{\alpha}$ can be chosen to be a vector in the null space of $E_{\alpha}$, which is the effective environment of $A_{\alpha}$ in $\langle\Psi_{0}|\Psi_{0}\rangle$ and satisfies $\langle\Psi_{0}|\Psi_{0}\rangle=\mathrm{Tr}E_{\alpha}A_{\alpha}$. For a given ground state with bond dimension $D$, we have $12D^{4}-$ 1 choices of $B_{\alpha}$ satisfying Eq.(S-3) for each $\alpha$, and thus $48D^{4}-$ 4 basis vectors, denoted as $|\Phi_{i}(k)\rangle$, in total to span the tangent space of the ground state. Then, we can calculate the matrix representation of the Hamiltonian in this non-orthogonal basis, and solving the generalized eigenvalue equation
$\langle\Psi_{i}(k)|\hat{H}|\Psi_{j}(k)\rangle=E(k)\langle\Psi_{i}(k)|\Psi_{j}(k)\rangle$ (S-4)
gives the excitation energies $E(k)$s and excited eigenstates $|\Phi_{i}(k)\rangle$s at momentum $k$ In this work, we have employed an effective derivative trick and a regularization procedure, introduced in Ref. [63, 64], to determine the matrices that appear in Eq. (S-4).
@@
[…]
@@
where $w^{\alpha}_{n}(k)\equiv|\langle\Phi_{n}(k)|S_{k}^{\alpha}|\Psi_{0}\rangle|^{2}$ is the spectral weight for each mode. If regarded as the overlap of two excited-state wave functions, the spectral weight can be tackled similarly by the derivative trick [64]. In practical calculations, a Lorentzian broadening is employed to deal with the delta function and mimic the finite temperature effect.
II Linear spin wave and bond-operator theories for the SS model
To better understand the TN results in the antiferromagnetic (AFM) and the plaquette phases, we study the spin excitations in these two phases by using the linear spin wave (LSW) and bond-operator theories, respectively.
II.1 Linear spin wave theory for spin excitations in the antiferromagnetic phase
@@
$\mathbf{n}_{l}\cdot\mathbf{S}_{{\mathbf{r}}l}$ $=$ $S-b_{{\mathbf{r}}l}^{\dagger}b_{{\mathbf{r}}l},$ (S-6)
$(\mathbf{u}_{l}+i\mathbf{v}_{l})\cdot\mathbf{S}_{{\mathbf{r}}l}$ $=$ $({2S-b_{{\mathbf{r}}l}^{\dagger}b_{{\mathbf{r}}l}})^{\frac{1}{2}}b_{{\mathbf{r}}l},$ (S-7)
$(\mathbf{u}_{l}-i\mathbf{v}_{l})\cdot\mathbf{S}_{{\mathbf{r}}l}$ $=$ $b_{{\mathbf{r}}l}^{\dagger}({2S-b_{{\mathbf{r}}l}^{\dagger}b_{{\mathbf{r}}l}})^{\frac{1}{2}}.$ (S-8)
@@
$b_{{\mathbf{r}}l}=\sqrt{\frac{2}{N}}\sum_{\mathbf{k}\in{\text{MBZ}}}b_{\mathbf{k}l}e^{i\mathbf{R}_{{\mathbf{r}}l}\cdot\mathbf{k}}$ (S-9)
@@
$b_{{\mathbf{r}}l^{\prime}}=\sqrt{\frac{2}{N}}\sum_{\mathbf{k}\in{\text{MBZ}}}b_{\mathbf{k}l^{\prime}}e^{-i\mathbf{R}_{{\mathbf{r}}l^{\prime}}\cdot\mathbf{k}}$ (S-10)
@@
$H_{\text{LSW}}=\sum_{\mathbf{k}\in{\text{MBZ}}}\Psi(\mathbf{k}){}^{\dagger}h(\mathbf{k})\Psi(\mathbf{k})+const.,$ (S-11)
@@
$\Psi(\mathbf{k})=[b_{\mathbf{k}1},b_{\mathbf{k}2},b^{\dagger}_{\mathbf{k}3},b^{\dagger}_{\mathbf{\mathbf{k}}4}]^{T},$ (S-12)
and
$h(k)=\left(\begin{array}[]{cccc}2J-\frac{J^{\prime}}{2}&\frac{J^{\prime}}{2}e^{\frac{i(k_{x}-k_{y})}{2}}&J\cos(\frac{k_{x}}{2})&J\cos(\frac{k_{y}}{2})\\ \frac{J^{\prime}}{2}e^{\frac{-i(k_{x}-k_{y})}{2}}&2J-\frac{J^{\prime}}{2}&J\cos(\frac{k_{y}}{2})&J\cos(\frac{k_{x}}{2})\\ J\cos(\frac{k_{x}}{2})&J\cos(\frac{k_{y}}{2})&2J-\frac{J^{\prime}}{2}&\frac{J^{\prime}}{2}e^{\frac{-i(k_{x}+k_{y})}{2}}\\ J\cos(\frac{k_{y}}{2})&J\cos(\frac{k_{x}}{2})&\frac{J^{\prime}}{2}e^{\frac{i(k_{x}+k_{y})}{2}}&2J-\frac{J^{\prime}}{2}\end{array}\right).$ (S-13)
@@
$\Phi(\mathbf{k})=[\beta_{\mathbf{k}1},\beta_{\mathbf{k}2},\beta_{\mathbf{k}3}^{\dagger},\beta_{\mathbf{k}4}^{\dagger}]^{T},$ (S-14)
@@
$\begin{matrix}E_{+}&=\sqrt{\frac{2J^{2}(J-J^{\prime})}{2J-J^{\prime}}}k_{x}+\frac{J^{2}J^{\prime}}{2(2J-J^{\prime})^{2}}k_{x}^{2},\\ \\ E_{-}&=\sqrt{\frac{2J^{2}(J-J^{\prime})}{2J-J^{\prime}}}k_{x}-\frac{J^{2}J^{\prime}}{2(2J-J^{\prime})^{2}}k_{x}^{2},\end{matrix}$ (S-15)
@@
$\begin{matrix}E_{+}&=\frac{Jk_{x}^{2}}{2},\\ \\ E_{-}&=0.\end{matrix}$ (S-16)
II.2 Bond operator theory for the triplet excitations in the plaquette phase
@@
Define a vacuum $|0\rangle$ state and four boson operators which yield the four physical states (one singlet and three triplets) by $|s\rangle=s^{\dagger}|0\rangle$, $|t_{\alpha}\rangle=t_{\alpha}^{\dagger}|0\rangle$, respectively [78, 77]. The projection operators are expressed $Z^{st_{\alpha}}=s^{\dagger}t_{\alpha}$, $Z^{t_{\alpha}t_{\beta}}=t_{\alpha}^{\dagger}t_{\beta}$. The spins represented via these boson operators are
$S_{l}^{\alpha}=\frac{(-1)^{l}}{\sqrt{6}}(s^{\dagger}t_{\alpha}+t_{\alpha}^{\dagger}s)-\frac{i}{4}\sum_{\beta,\gamma}\epsilon_{\alpha\beta\gamma}t_{\beta}^{\dagger}t_{\gamma}$ (S-17)
@@
$H=H_{0}+\sum_{\mathbf{k}\in{\text{MBZ}}}\Psi(\mathbf{k}){}^{\dagger}h(\mathbf{k})\Psi(\mathbf{k})+const.,$ (S-18)
@@
$\Psi(\mathbf{k})=[t_{\mathbf{k}x},t_{\mathbf{k}y},t_{\mathbf{\mathbf{k}}z},t^{\dagger}_{\mathbf{-k}x},t^{\dagger}_{\mathbf{-k}y},t^{\dagger}_{\mathbf{\mathbf{-k}}z}]^{T}$ (S-19)
@@
Then following the stantard Bogoliubov diagonalization we can derive the dispersion of the lowest triplet as
$E=\sqrt{J^{2}+\frac{2}{3}J(J^{\prime}-2J)(\cos(k_{x})+\cos(k_{y}))}.$ (S-20)
@@
We compare the triplet dispersion with that calculated by using the TN method in Fig. 3 of the main text, and they agree well. The small difference is likely caused by the finite bond dimension $D$ in the TN calculation.
III Numerical results on the spin excitation spectra

### v2

@@ Acknowledgements.
We thank Runze Chi, Wenan Guo, Bruce Normand, Wei Li, Zhengxin Liu, Anders W. Sandvik, Yiming Wang, Jize Zhao for helpful discussions. This work is supported by the National Key R&D Program of China (Grant No. 2023YFA1406500), the National Natural Science Foundation of China (Grant Nos. 12334008, 12274458, 12174441).
@@
Appendix A iPEPS method for calculating the ground state and excited states
@@
Figure 5: (a): The iPEPS setup in the calculation of the ground state of the SS model. The tensors are arranged at the sites of the SS lattice. The bright-colored sites indicate a $2\times 2$ unit cell used in the calculation. (b): Illustration of the single-mode iPEPS ansatz for the calculation of excited states. In each term of the summation, a single local tensor $A_{\alpha}$ located at site $r$ belonging to a specified sublattice is modified to a different tensor $B_{\alpha}$. Here $\alpha=1,2,3,4$, corresponding to a $2\times 2$ unit cell.
@@
In this work, we apply the infinite projected entangled pair state (iPEPS) method verstraete2004renormalization ; orus2014practical to calculate the ground state of the Shastry-Sutherland (SS) model. A sketched iPEPS setup with a $2\times 2$ unit cell is shown in Fig. 5(a). In the iPEPS setup, a local tensor is defined on each site of the SS lattice. As shown in Fig.  5, each local tensor has a physical index corresponding to the spin degrees of freedom, with dimension $2$ and denoted as curve, and four virtual indices corresponding to the four links of the site, with dimension $D$ referred to as the iPEPS bond dimension and denoted as straight. In this work, as emphasized as bright color in Fig. 5, the unit cell is chosen to be $2\times 2$, and this is the minimal size compatible with the ground state of the SS model.
@@
$E$ $=$ $\frac{\sum_{\langle ij\rangle}\langle\Psi|\hat{H}_{ij}(J)|\Psi\rangle}{\langle\Psi|\Psi\rangle}$ (4)
$+$ $\frac{\sum_{\langle\langle ij\rangle\rangle}\langle\Psi|\hat{H}_{ij}(J^{\prime})|\Psi\rangle}{\langle\Psi|\Psi\rangle}$
where $\langle ij\rangle$ and $\langle\langle ij\rangle\rangle$ denote nearest neighbors and second nearest neighbors, respectively. In this sense, the energy can be regarded as a complicated function of the local tensors, the variational parameters in the iPEPS ansatz. The denominators and numerators in Eq. (4) can be effectively calculated by using the corner transfer matrix renormalization group method corboz2014competing ; orus2009simulation ; PhysRevB.84.041108 ; VCTMRG2022 . Then, starting from an arbitrary state, the ground state is obtained by variationally minimizing the energy through gradient-based optimization, e.g., the L-BFGS optimization strategy LBFGS . In this work, the energy gradients concerning the local tensors are obtained effectively by the automatic differentiation techniques SRGAD ; liao2019differentiable .
Once the ground state $|\Psi_{0}\rangle$ is obtained, we can construct the iPEPS ansatz for the excited state through the single-mode approximation ponsioen2020excitations ; ponsioen2022automatic ; chi2022spin . More specifically, the excited ansatz with lattice momentum $k$ reads as
$|\Psi_{k}(B_{\alpha})\rangle=\sum_{r\in\alpha}e^{ikr}|\Psi_{r}(B_{\alpha})\rangle$ (5)
@@
$\langle\Psi_{r}(B_{\alpha})|\Psi_{0}\rangle=0.$ (6)
In practice, $B_{\alpha}$ can be chosen to be a vector in the null space of $E_{\alpha}$, which is the effective environment of $A_{\alpha}$ in $\langle\Psi_{0}|\Psi_{0}\rangle$ and satisfies $\langle\Psi_{0}|\Psi_{0}\rangle=\mathrm{Tr}E_{\alpha}A_{\alpha}$. For a given ground state with bond dimension $D$, we have $2D^{4}-1$ choices of $B_{\alpha}$ satisfying Eq.(6) for each $\alpha$, and thus $8D^{4}-4$ basis vectors, denoted as $|\Phi_{i}(k)\rangle$, in total to span the tangent space of the ground state. Then, we can calculate the matrix representation of the Hamiltonian in this non-orthogonal basis, and solving the generalized eigenvalue equation
$\langle\Psi_{i}(k)|\hat{H}|\Psi_{j}(k)\rangle=E(k)\langle\Psi_{i}(k)|\Psi_{j}(k)\rangle$ (7)
gives the excitation energies $E(k)$s and excited eigenstates $|\Phi_{i}(k)\rangles$ at momentum $k$ . In this work, we have employed an effective derivative trick and a regularization procedure, introduced in Ref. ponsioen2020excitations ; ponsioen2022automatic , to determine the matrices that appear in Eq. (7).
@@
$\mathcal{S}^{\alpha\alpha}(k,\omega)$ $=$ $\int dte^{-i\omega t}\langle\Psi_{0}|S_{-k}^{\alpha}S^{\alpha}_{k}(t)|\Psi_{0}\rangle$ (8)
$=$ $\sum_{mn}\int dte^{-i(\omega+E_{0})t}\langle\Psi_{0}|S_{-k}^{\alpha}|\Phi_{m}(k)\rangle$
$\cdot$ $\langle\Phi_{m}(k)|e^{iHt}|\Phi_{n}(k)\rangle\langle\Phi_{n}(k)|S^{\alpha}_{k}|\Psi_{0}\rangle$
@@
where $w^{\alpha}_{n}(k)\equiv|\langle\Phi_{n}(k)|S_{k}^{\alpha}|\Psi_{0}\rangle|^{2}$ is the spectral weight for each mode. If regarded as the overlap of two excited-state wave functions, the spectral weight can be tackled similarly by the derivative trick ponsioen2022automatic . In practical calculations, a Lorentzian broadening is employed to deal with the delta function and mimic the finite temperature effect.
Appendix B Linear spin wave and bond-operator theories for the SS model
To better understand the tensor network results in the antiferromagnetic (AFM) and the plaquette phases, we study the spin excitations in these two phases by using the linear spin wave (LSW) and bond-operator theories, respectively.
B.1 Linear spin wave theory for spin excitations in the antiferromagnetic phase
@@
$\mathbf{n}_{l}\cdot\mathbf{S}_{{\mathbf{r}}l}$ $=$ $S-b_{{\mathbf{r}}l}^{\dagger}b_{{\mathbf{r}}l},$ (9)
$(\mathbf{u}_{l}+i\mathbf{v}_{l})\cdot\mathbf{S}_{{\mathbf{r}}l}$ $=$ $({2S-b_{{\mathbf{r}}l}^{\dagger}b_{{\mathbf{r}}l}})^{\frac{1}{2}}b_{{\mathbf{r}}l},$ (10)
$(\mathbf{u}_{l}-i\mathbf{v}_{l})\cdot\mathbf{S}_{{\mathbf{r}}l}$ $=$ $b_{{\mathbf{r}}l}^{\dagger}({2S-b_{{\mathbf{r}}l}^{\dagger}b_{{\mathbf{r}}l}})^{\frac{1}{2}}.$ (11)
@@
$b_{{\mathbf{r}}l}=\sqrt{\frac{2}{N}}\sum_{\mathbf{k}\in{\text{MBZ}}}b_{\mathbf{k}l}e^{i\mathbf{R}_{{\mathbf{r}}l}\cdot\mathbf{k}}$ (12)
@@
$b_{{\mathbf{r}}l^{\prime}}=\sqrt{\frac{2}{N}}\sum_{\mathbf{k}\in{\text{MBZ}}}b_{\mathbf{k}l^{\prime}}e^{-i\mathbf{R}_{{\mathbf{r}}l^{\prime}}\cdot\mathbf{k}}$ (13)
@@
$H_{\text{LSW}}=\sum_{\mathbf{k}\in{\text{MBZ}}}\Psi(\mathbf{k}){}^{\dagger}h(\mathbf{k})\Psi(\mathbf{k})+const.,$ (14)
@@
$\Psi(\mathbf{k})=[b_{\mathbf{k}1},b_{\mathbf{k}2},b^{\dagger}_{\mathbf{k}3},b^{\dagger}_{\mathbf{\mathbf{k}}4}]^{T},$ (15)
and $h(k)=$
$\left(\begin{array}[]{cccc}2J-\frac{J^{\prime}}{2}&\frac{J^{\prime}}{2}e^{\frac{i(k_{x}-k_{y})}{2}}&J\cos(\frac{k_{x}}{2})&J\cos(\frac{k_{y}}{2})\\ \frac{J^{\prime}}{2}e^{\frac{-i(k_{x}-k_{y})}{2}}&2J-\frac{J^{\prime}}{2}&J\cos(\frac{k_{y}}{2})&J\cos(\frac{k_{x}}{2})\\ J\cos(\frac{k_{x}}{2})&J\cos(\frac{k_{y}}{2})&2J-\frac{J^{\prime}}{2}&\frac{J^{\prime}}{2}e^{\frac{-i(k_{x}+k_{y})}{2}}\\ J\cos(\frac{k_{y}}{2})&J\cos(\frac{k_{x}}{2})&\frac{J^{\prime}}{2}e^{\frac{i(k_{x}+k_{y})}{2}}&2J-\frac{J^{\prime}}{2}\end{array}\right).$ (16)
@@
$\Phi(\mathbf{k})=[\beta_{\mathbf{k}1},\beta_{\mathbf{k}2},\beta_{\mathbf{k}3}^{\dagger},\beta_{\mathbf{k}4}^{\dagger}]^{T},$ (17)
@@
$\begin{matrix}E_{+}&=\sqrt{\frac{2J^{2}(J-J^{\prime})}{2J-J^{\prime}}}k_{x}+\frac{J^{2}J^{\prime}}{2(2J-J^{\prime})^{2}}k_{x}^{2},\\ \\ E_{-}&=\sqrt{\frac{2J^{2}(J-J^{\prime})}{2J-J^{\prime}}}k_{x}-\frac{J^{2}J^{\prime}}{2(2J-J^{\prime})^{2}}k_{x}^{2},\end{matrix}$ (18)
@@
$\begin{matrix}E_{+}&=\frac{Jk_{x}^{2}}{2},\\ \\ E_{-}&=0.\end{matrix}$ (19)
B.2 Bond operator theory for the triplet excitations in the plaquette phase
@@
Define a vacuum $|0\rangle$ state and four boson operators which yield the four physical states (one singlet and three triplets) by $|s\rangle=s^{\dagger}|0\rangle$, $|t_{\alpha}\rangle=t_{\alpha}^{\dagger}|0\rangle$, respectively sachdev1990bond ; zhitomirsky1996valence . The projection operators are expressed $Z^{st_{\alpha}}=s^{\dagger}t_{\alpha}$, $Z^{t_{\alpha}t_{\beta}}=t_{\alpha}^{\dagger}t_{\beta}$. The spins represented via these boson operators are
$S_{l}^{\alpha}=\frac{(-1)^{l}}{\sqrt{6}}(s^{\dagger}t_{\alpha}+t_{\alpha}^{\dagger}s)-\frac{i}{4}\sum_{\beta,\gamma}\epsilon_{\alpha\beta\gamma}t_{\beta}^{\dagger}t_{\gamma}$ (20)
@@
$H=H_{0}+\sum_{\mathbf{k}\in{\text{MBZ}}}\Psi(\mathbf{k}){}^{\dagger}h(\mathbf{k})\Psi(\mathbf{k})+const.,$ (21)
@@
$\Psi(\mathbf{k})=[t_{\mathbf{k}x},t_{\mathbf{k}y},t_{\mathbf{\mathbf{k}}z},t^{\dagger}_{\mathbf{-k}x},t^{\dagger}_{\mathbf{-k}y},t^{\dagger}_{\mathbf{\mathbf{-k}}z}]^{T}$ (22)
@@
Then following the standard Bogoliubov diagonalization we can derive the dispersion of the lowest triplet as
$E=\sqrt{J^{2}+\frac{2}{3}J(J^{\prime}-2J)(\cos(k_{x})+\cos(k_{y}))}.$ (23)
@@
We compare the triplet dispersion with that calculated by using the tensor network method in Fig. 3 of the main text, and they agree well. The small difference is likely caused by the finite bond dimension $D$ in the tensor network calculation.
Appendix C Numerical results on the spin excitation spectra